OpenStudy (anonymous):

Trigonometry? what is the smallest value for x where y =sin2x reaches its maximum? explain please

OpenStudy (anonymous):

sine is biggest when it is 1, and \(\sin(\frac{\pi}{2})=1\) so set \[2x=\frac{\pi}{2}\] and solve for \(x\) in one step

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

yw